OpenStudy (moongazer):

The equation of a circle center at the origin and radius 1is ?

OpenStudy (anonymous):

x^2 +y^2 =1

OpenStudy (moongazer):

explain why please ill give you a medal if you explained it well

OpenStudy (anonymous):

there isnt anyway of explaining it really lol

OpenStudy (anonymous):

the general form a circle with centre (a,b) and radius r is \[(x-a)^2 +(y-b)^2 =r^2\]

OpenStudy (anonymous):

here centre is (0,0) yeh.. theres not alot of explaining, you have to remember the formula , thats all you can do

OpenStudy (anonymous):

i can explain

OpenStudy (anonymous):

ok lets go back to geometry definition

OpenStudy (anonymous):

you there?

OpenStudy (moongazer):

ok

OpenStudy (anonymous):

ohh its phytagorus , yeh

OpenStudy (moongazer):

yes

OpenStudy (anonymous):

the locus definition

OpenStudy (anonymous):

a circle is a set of points such that, the distance from the center is fixed

OpenStudy (anonymous):

right

OpenStudy (anonymous):

a circle is a locus of points whose distance from a point is constant, that distance is the radius

OpenStudy (anonymous):

thats the ancient greek definition, going back when

OpenStudy (anonymous):

now fast forward to cartesian world, we have x y graph

OpenStudy (anonymous):

your center lets say for convenience is origin (0,0)

OpenStudy (anonymous):

ok so far?

OpenStudy (anonymous):

your doing it the really long way, not very good explaination lol

OpenStudy (anonymous):

ill do it

OpenStudy (moongazer):

ok but please explain easier to understand

OpenStudy (anonymous):

he wants a thorough explanation

OpenStudy (anonymous):

we want d( (x,y) (0,0) = r

OpenStudy (anonymous):

the distance between (0,0) and the points (x,y) to be the constant radius

OpenStudy (anonymous):

ok so far? , and we know the distance between any two points using the distance formula

OpenStudy (anonymous):

elec, this is not easy to understand? i can show a graph , but this interface sucks balls

OpenStudy (anonymous):

you can just go straight to the general case centre (a,b) \[\sqrt{(x-a)^2 +(y-b)^2} = r \]

OpenStudy (moongazer):

please explain it like you are explaining it to a kid

OpenStudy (anonymous):

then square both sides

OpenStudy (anonymous):

d ( (x1,y1) (x2,y2) ) = sqrt ( (x1-x2)^2 + (y1-y2)^2)

OpenStudy (anonymous):

right that works

OpenStudy (moongazer):

WAIT!

OpenStudy (anonymous):

so the distance between any arbitrary center (h,k) and the points (x,y) is some distance r. d ( h,k)(x,y) = r

OpenStudy (anonymous):

, so sqrt ( (x-h)^2 + ( y - k)^2) = r

OpenStudy (anonymous):

this is a deductive proof my friend :)

OpenStudy (moongazer):

my teacher gave me a really easy solution for this and i just forgot it and then i wrote the equation x^2+y^2=r^2 and i forgot what it means

OpenStudy (moongazer):

on my ntbk

OpenStudy (anonymous):

if you draw a line from your center to a point on the circle, and then draw a triangle

OpenStudy (anonymous):

you get pythagorean triangle

OpenStudy (moongazer):

please don't give me a complicated solution

OpenStudy (anonymous):

ok youre annoying, bye

OpenStudy (moongazer):

then?

OpenStudy (moongazer):

youre doing it right

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

lols

OpenStudy (anonymous):

then the world ends tomorrow

OpenStudy (anonymous):

brb, i will use a whiteboard , this is no good

OpenStudy (moongazer):

ahhhh ok

OpenStudy (anonymous):

a more interesting question has peaked my curiosity, brb

OpenStudy (moongazer):

ill just give you both a medal^_^

OpenStudy (moongazer):

for answering and explaining^_^

OpenStudy (anonymous):

complex numbers: find a, b such that (-bi)^2 = a^2 ?

OpenStudy (anonymous):

this question is ambiguous, ok one sec

OpenStudy (anonymous):

here click on this http://www.twiddla.com/542364

OpenStudy (moongazer):

what is this?

OpenStudy (anonymous):

circle centre (2,3), radius 2

OpenStudy (anonymous):

its pretty basic

OpenStudy (moongazer):

it is alright know^_^